The boiling point of chloroform was raised by 0.30K when 5×10⁻⁴ kg of an organic compound was dissolved in 3×10⁻² kg of chloroform. To calculate the molar mass of organic compound. Molal elevation constant for chloroform is 3.9K kg mol⁻¹.
1
Expert's answer
2020-12-21T03:53:58-0500

The boiling point elevation can be calculated using the following equation:

"\Delta T = K_bm" ,

where "\Delta T" is the change in the temperature, "K_b" is the molal boiling point elevation constant and "m" is the molal concentration of the solute in the solution.

From this equation, the molal concentration of the organic compound is:

"m = \frac{\Delta T}{K_b} = \frac{0.30}{3.9} = 0.0769" mol/kg.

The molality, or the molal concentration is defined as the number of the moles of the solute divided by the mass of the solvent (chloroform is our case):

"m = \frac{n}{m_{solvent}}" .

Therefore, the number of the moles of the solute present in the solution is:

"n = m\u00b7m_{solvent} = 0.0769\u00b70.03 = 0.00231" mol.

Finally, the molar mass of the solute is:

"M = \frac{m_{solute}}{M} = \frac{5·10^{-4}·1000(g)}{0.00231(mol)} = 216.7" g/mol.


Answer: the molar mass of the organic compound dissolved in chloroform is 216.7 g/mol.